# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v2

@@ VI-A Results on NTU4DRadLM
We compare our odometry with previous published results on the NTU4DRadLM dataset [29], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [30], as well as other works such as EFEAR-4D [26] or RIV-SLAM [24]. We include these self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public. Even though there are no published results for EKF-RIO [4] on this dataset, it is one of the few methods with available source code, and so we are able to produce and include evaluation results for it. We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
@@
Table I shows the comparison between all methods. We can observe how our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation). Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
@@ VI-B Results on Snail-Radar
In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [8], which uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions. In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform. These sequences are all fairly long, with the longest (iaf_20231213_2) being 15 min and most being around 7 min in length.
Table II shows results using the Oculii Eagle radar. In general, we can observe the best results to be spread across the different methods. Nonetheless, our Gaussian algorithms obtain the best or second best results in 9 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT). The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences. It is worth mentioning that NDT, unlike in NTU4DRadLM, seems to behave much better in Snail-Radar.
Table III shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be caused by the radar lacking the scan enhancement algorithm used by Oculii Eagle. However interestingly, we can observe better metrics in general compared to Oculii Eagle. The sparser scans trigger a known issue in PCL’s implementation of NDT that cause the program to abnormally terminate, so unfortunately NDT is excluded from the comparison. The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results. Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version. We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
@@ VI-C Summary of results
We consider that the results described above validate our methodology. Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments. Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results. Our EKF-based radar-inertial framework is also validated by the experiments, which show how on a publicly evaluated dataset we are able to improve on the state of the art even when not using our Gaussian algorithms. On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
@@ VII Conclusion
In this paper we introduce a Gaussian-based radar-inertial odometry pipeline. Our methods summarize noisy 4D radar scans using freely positioned 3D Gaussians (like 3DGS [10] and unlike NDT [2]), which then support multi-hypothesis scan matching for increased robustness. Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences.
Future works include improving the Gaussian models with an optimization algorithm more efficient than gradient descent, better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as intensity values; improving the robustness of the multi-hypothesis particle system with kinematic priors, dynamically adjusting the hyperparameters of the system, improving the quality of egovelocity estimation by modeling kinematic restrictions relevant to the robot platform used, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.

### v3

@@ VI-A Results on NTU4DRadLM
We evaluate using the NTU4DRadLM dataset [8], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU. Besides the evaluation performed with our pipeline, we also include previously published results of several radar-focused odometry methods, in order to provide context on the current performance of the state of the art. Due to a lack of Radar-Inertial methods, we also include radar-only methods. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [9], as well as other works such as EFEAR-4D [10] or RIV-SLAM [11]. All works are categorized according to the sensory input used (IMU, egovelocity, scan matching). We include self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public. The results for EKF-RIO [29] are generated by running the system directly, as there are no published results on this dataset. We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
@@
Table I shows the comparison between all methods. Our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation). Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
Additionally, we report the average scan matching processing time for each method running under our odometry pipeline. The evaluation platform is a desktop PC equipped with an Intel Core i9-9900X CPU and 128 GB of RAM. The single-hypothesis version of our method achieves the fastest scan matching processing times, closely followed by VGICP. The multi-hypothesis approach, despite simultaneously registering 8 particles at once, is only 3.62x slower than the single-hypothesis version. We also mention that our current RIO implementation is entirely sequential, and that Gaussian modeling at keyframes consumes up to 608.78ms, raising the average scan processing time to 18.13ms. However, we believe that running the modeling and odometry tasks in parallel would remove any timing inconsistencies affecting real-time usage; which we leave as future work.
@@ VI-B Results on Snail-Radar
In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [31], which uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions. In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform. The longest tested sequence (iaf_20231213_2) is 15 min in length, while the others are around 7 min.
Table II shows results using the Oculii Eagle radar. In general, we can observe the best results to be spread across the different methods. Nonetheless, our Gaussian algorithms obtain the best or second best results in 9 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT). The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences. We also note that NDT seems to behave much better in Snail-Radar than in NTU4DRadLM.
Table III shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be the effect of Oculii Eagle’s scan enhancement algorithm. However, we can observe better metrics in general compared to Oculii Eagle. The sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from the comparison. The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results. Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version. We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
@@ VI-C Summary of results
We consider that the results above validate our methodology. Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments. Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results. On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
@@ VII Conclusion
In this paper we introduce modeling and scan matching algorithms based on 3D Gaussians, which initialize and jointly optimize all parameters of a 3D Gaussian model to fit a radar point cloud, and make use of multiple hypothesis to support a scan matching process for subsequent incoming point clouds. Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences. At the same time, our methods offer higher flexibility/expandability and faster scan matching compared to existing approaches based on individual points and voxelization [2, 1, 15, 16].
Future works include improving the run-time efficiency of the Gaussian modeling (replacing gradient descent, and running in parallel to the odometry pipeline), better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as RCS (Radar Cross Section) values, and improving the robustness of the multi-hypothesis particle system with kinematic priors. Another line of work is further development of the RIO pipeline, using techniques such as Pose Graph Optimization (PGO), loop closure, and creating a global Gaussian map of the environment for the purposes of full SLAM.